OpenStudy (anonymous):

find the slope and the y intercept: f(x)=-8x-3

OpenStudy (anonymous):

y=mx+b m is slope b is y intercept what's in the place of m? what's in the place of b

OpenStudy (anonymous):

slope = -8/1 y intercept = -3

OpenStudy (anonymous):

your given is already your answer.. xD